OpenStudy (anonymous):

sin h 3x e^4x find the nth derivative

OpenStudy (anonymous):

Is that: \[sinh{3x} \cdot e^{4x}\] ?

OpenStudy (anonymous):

what to do with that?

OpenStudy (anonymous):

if the eq is like traxter wrote, the exchange sinh(3x) as a exp function.

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

sin h 3x e^4x find the nth derivative

hartnn (hartnn):

can u find the n th derivative of e^{ax} ?

OpenStudy (anonymous):

sin h 3x e^4x find the nth derivative will u derive it

OpenStudy (anonymous):

can u find the n th derivative of e^{ax} ? form:a^n e^ax

OpenStudy (anonymous):

e^ax nth derivative is (a^n)e^ax

OpenStudy (anonymous):

maybe it is easy to change sinh3x in a exp function

OpenStudy (anonymous):

Leibniz' theorem/ binomial stuff will work for u\[(uv)^{(n)}=\sum_{k=0}^{n}\left(\begin{matrix}n\\k\end{matrix}\right)u^{(n-k)}v^k\]

OpenStudy (anonymous):

sinh3x = (e^3x - e^-3x)/2

OpenStudy (anonymous):

so the given eq is ruduced as (e^7x - e^x)/2

OpenStudy (anonymous):

then you apply a e^ax derivative rule

OpenStudy (anonymous):

can u explain breifly plz

OpenStudy (anonymous):

then the answer is [(7^n)e^7x - e^x]/2

OpenStudy (anonymous):

the definition of the sinhx = (e^x - e^-x)/2. we promised it

OpenStudy (anonymous):

ignore my comment

hartnn (hartnn):

@shambavi ask if u still don't get it.

OpenStudy (anonymous):

@LeeYeongKyu heyy thankxxxx i got it

OpenStudy (anonymous):

if you want to know why sinhx = (e^x - e^-x)/2 use euler theory